A rug is to fit in a room so that a border of even width is left on all four sides. If the room is 12 feet by 15 feet and the area of the rug is 108 square feet, how wide will the border be?
Answer by Boreal(15235)

(12-2x)(15-2x) is the rug fitting in the room, with borders equal and are x.
This expands to 180-54x+4x^2=108, units in feet, product in square feet
4x^2-54x-72=0
(2x-3)(2x-24)=0
x=12 (won't work), 1.5
The border is 1.5 feet wide.
The rug is 9 *12 feet or 108 square feet.